OpenStudy (anonymous):

Using the following equation, find the center and radius of the circle. You must show all work and calculations to receive credit. x2 + 2x + y2 + 4y = 19

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

I've gotten this far idk what to do next x2 + 2x + y2 + 4y = 19 (x2 + 2x) + (y2 + 4y) = 19

ganeshie8 (ganeshie8):

good progress, next is easy just complete the square for each bracket

OpenStudy (anonymous):

how do you do that

ganeshie8 (ganeshie8):

do u knw (a+b)^2 formula ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

sorry

ganeshie8 (ganeshie8):

its okay :) here is the formula :- (a + b)^2 = a^2 + 2ab + b^2

ganeshie8 (ganeshie8):

basically, we want to cookup the brackets into something like above

ganeshie8 (ganeshie8):

x2 + 2x + y2 + 4y = 19 (x2 + 2x) + (y2 + 4y) = 19 (x2 + 2*1*x) + (y2 + 2*2*y) = 19

ganeshie8 (ganeshie8):

fine wid above ?

OpenStudy (anonymous):

so do you mulity them together to get the next step

ganeshie8 (ganeshie8):

yes, i just pulled the 2 out

ganeshie8 (ganeshie8):

next ima add 1^2 in first brac, and 2^2 in second brac

OpenStudy (anonymous):

so itll be (x2 + 2x+2) + (y2 + 4y+ 4) = 19

ganeshie8 (ganeshie8):

x2 + 2x + y2 + 4y = 19 (x2 + 2x) + (y2 + 4y) = 19 (x2 + 2*1*x) + (y2 + 2*2*y) = 19 (x^2 + 2*1*x + 1^2) + (y^2 + 2*2*y + 2^2) = 19 + 1^2 + 2^2 (x+1)^2 + (y+2)^2 = 24

ganeshie8 (ganeshie8):

ive used (a+b)^2 formula in last step, above

ganeshie8 (ganeshie8):

let me knw once u digest above

OpenStudy (anonymous):

so is that the final answer??? @ganeshie8

ganeshie8 (ganeshie8):

humm ased on above equation, center = (-1, -2) radius = sqrt(24)

OpenStudy (anonymous):

Can you please help me with another question?

OpenStudy (anonymous):

Stacy wants to construct a circle inscribed in a triangle while Ben wants to construct a circle circumscribed about a triangle. How will Stacy and Ben's steps differ from each other?

OpenStudy (anonymous):

@ganeshie8